A rectangular table has a perimeter of 18 feet. Its length is 5 feet greater than its width what is the demension
Answer by JulietG(1812)

Perimeter (think, a fence around) equals 2 width + 2 length
P = 2W + 2L
18 = 2W + 2L
Divide by 2
9 = W + L
L = W+5
Substitute the value of L into the equation above it.
9 = W+L
9 = W+(W+5)
9 = 2W+5
Subtract 5 from each side
4 = 2W
Divide each side by 2
2 = W
If the width is 2, then the length is 5 more than that, or 7
W = 2; L = 7
P = 2+2+7+7, or 18